OpenStudy (anonymous):

THEREV ARE APPLES,ORANGES AND PEARS IUN BETH REFRIGERATOR. SHE HAS 3 APPLES AND 2 ORANGES .IN ALL , SHE HAS 11 PIECES OF FRUIT . HOW MANY PEARS DOES SHE HAVE.

myininaya (myininaya):

let apples be represented by a oranges by o pears by p and fruit in all represent f a+o+p=f we know that we have a=3 and o=2 and f=11 plug does in and solve for p

myininaya (myininaya):

3+2+p=11

myininaya (myininaya):

3+2=?

OpenStudy (anonymous):

Why is this guy using all CAPS? It is bad on people's eyes.

OpenStudy (karatechopper):

lol

OpenStudy (karatechopper):

can i answer the 3+2?

myininaya (myininaya):

maybe he accidently pressed the caps lock button you can only answer it if you can get it right chopper lol